Six Measures of Solvency or Profitability

The following data were taken from the financial statements of Gates Inc. for the current fiscal year.

Property, plant, and equipment (net) $1,162,000
Liabilities:
Current liabilities $166,000
Note payable, 6%, due in 15 years 830,000
Total liabilities $996,000
Stockholders' equity:
Preferred $4 stock, $100 par (no change during year) $747,000
Common stock, $10 par (no change during year) 747,000
Retained earnings:
Balance, beginning of year $796,000
Net income 291,000 $1,087,000
Preferred dividends $29,880
Common dividends 61,120 91,000
Balance, end of year 996,000
Total stockholders' equity $2,490,000
Sales $21,073,800
Interest expense $49,800

Assuming that total assets were $3,312,000 at the beginning of the current fiscal year, determine the following. When required, round to one decimal place.

a. Ratio of fixed assets to long-term liabilities
b. Ratio of liabilities to stockholders' equity
c. Asset turnover
d. Return on total assets %
e. Return on stockholders’ equity %
f. Return on common stockholders' equity %
Answer #1

a. Ratio of fixed assets to long-term liabilities = Fixed assets / Long term Liabilities
= $1162000/830000 = 1.4 : 1

b. Ratio of liabilities to stockholders' equity = Total Liabilities / Total Stockholders Equity
= $996000 / 2490000 = 0.4

c. Asset Turnover = Sales / Average Total Assets
= $21073800 / 3399000 = 6.2
Average Total Assets = ($996000+2490000+3312000)/2 =$3399000

d. Return on total Assets = Net Income / Average Total Assets
= $291000/3399000 = 8.56%

alternatively = 291000/3312000 = 8.79%

e. Return on stockholders’ equity = Net Income / Total Stockholders Equity
= $291000 / 2490000 = 11.69%

f. Return on common stockholders' equity = Net Income for common stockholder / Total Common Stockholders Equity
= ($291000-29880)/(747000+996000) = 14.98%